# Does Writing with Language Models Reduce Content Diversity?

### v2

@@ 3 How much do users engage with the model?
While I believe the concerns regarding children’s screen time are valid, I believe it is somewhat biased to not take this problem, which is a genuine issue right now, as an everyone problem, [skipped] I know that I, along with many other teenagers, would like nothing more than to go back to school, play sports outside, meet new people, and such. [skipped] They are also places for new ideas, watching college lectures, and political discourse [skipped] The problem of screen time should be considered an everyone problem, not just a student one Social media can be used for educational purposes Limiting screen time may not be effective in the long run Parents should trust their teenager more and not worry too much about their screen time
@@ 5.2 Results
keeping up with the news
7
keep up with the news
14
in my opinion the most
7
on animal welfare when humans
12
keep up with the news
6
to focus on animal welfare
11
opinion the most important things
6
selfish to pursue risky sports
11
but on the other hand
5
students should learn in school
11
the most important thing that
5
wrong to focus on animal
10
wrong to focus on animal
5
sports like extreme mountain climbing
10
focus on animal welfare when
5
keeping up with the news
9
unfair when it is considered
5
the end of the day
9
in my opinion listening to
4
things students should learn in
9
@@ 7 Related Work
Traditional evaluation metrics primarily assess text similarity to reference text (Papineni et al., 2002; Lin, 2004; Sellam et al., 2020). In interactive contexts where references are absent, alternative metrics gauge model assistance effectiveness indirectly. These include user-rated model helpfulness (Roemmele and Gordon, 2018; Clark et al., 2018; Padmakumar and He, 2022), retention of model suggestions in the final text (Akoury et al., 2020), and completion time (Buschek et al., 2021). Lee et al. (2022b) additionally considers user ownership and enjoyment which emphasizes maintaining user writing style. Our work extends existing metrics that evaluate lexical diversity and text style based on token and POS n-grams statistics (Roemmele et al., 2017; See et al., 2019; Tevet and Berant, 2021; Meister et al., 2023). We adapt these metrics for co-writing scenarios and introduce content-specific diversity measures.
@@ Acknowledgements
We would like to thank Mina Lee, Nick Lourie, Naomi Saphra, Richard Pang, and Nitish Joshi for their input at various stages of the project. We would also like to thank Abby Rabinowitz and Alexander Landfair from the NYU Expository Writing Program for valuable discussions during the planning stages. We would finally like to acknowledge all the writers recruited for this project without whom this work would not have been possible. This work is supported by Open Philanthropy, AWS AI, and the National Science Foundation under Grant No. 1922658.
@@ A.2 Decoding Parameters on OpenAI API
What Are the Most Important Things Students Should Learn in School? In your opinion, what are the most important things students should learn in school? What is the most important thing you have learned in school? How has this knowledge affected your life? How do you think it will help your success in the future? (Link)
stereotype
@@
What Stereotypical Characters Make You Cringe? What stereotypical characters in books, movies or television shows make you cringe and why? Would you ever not watch or read something because of its offensive portrayal of someone? (Link)
audiobook
@@
Is Listening to a Book Just as Good as Reading It? Do you listen to audiobooks? What are the benefits, in your opinion, of listening instead of reading? Are there advantages to reading that cannot be gained by listening? Which method do you prefer? Why? (Link)
athletes
@@
Should College Athletes Be Paid? Do you think college athletes should be paid? Or is a college scholarship and other non-monetary perks like the opportunity to play in front of cheering fans enough? What possible difficulties or downsides might there be in providing monetary compensation to players? (Link)
extremesports
@@
Is It Selfish to Pursue Risky Sports Like Extreme Mountain Climbing? Some sports, like extreme mountain climbing, are dangerous. Since there are varying degrees of risk in most, if not all, sports (such as the possibility of concussions, broken bones and even death), how does one decide where the line might be drawn between what is reasonable and what is not? Are some sports simply too dangerous to be called a sport? (Link)
animal
@@
Is It Wrong to Focus on Animal Welfare When Humans Are Suffering? Would you be surprised to hear that a study found that research subjects were more upset by stories of a dog beaten by a baseball bat than of an adult similarly beaten? Or that other researchers found that if forced to choose, 40 percent of people would save their pet dog over a foreign tourist. Why do you think many people are more empathetic toward the suffering of animals than that of people? In your opinion, is it wrong to focus on animal welfare when humans are suffering? Why do you think so? (Link)
news
@@
Are We Being Bad Citizens If We Don’t Keep Up With the News? In your opinion, are we being bad citizens if we don’t keep up with the news? Do you think all people have some responsibility to know what is going on in the world? Does engaging with current events actually do anything at all? Why do you think the way you do? (Link)
mindfulness
@@
Should Schools Teach Mindfulness? Have you ever tried mindfulness or meditation, practices that focus on the present moment and being aware of your thoughts, feelings and bodily sensations? If so, what was it like for you? If not, does it sound like something you’d like to try? Do you think that such practices have a place in schools? Why or why not? (Link)
screen
@@
How Worried Should We Be About Screen Time During the Pandemic? The coronavirus pandemic ended the screen time debate: Screens won. We all now find ourselves on our screens for school, for work and for connecting with family and friends during this time of social distancing and increased isolation. But should we be worried about this excessive screen use right now? Or should we finally get over it and embrace the benefits of our digital devices? (Link)
dating
@@
How Do You Think Technology Affects Dating? Have you had any experience with dating? Have you ever used dating apps? If so, what has it been like for you? If not, why not? How do you think technology — like apps, Netflix, social media and texting — affects dating and relationships? In your opinion, does it improve or worsen romantic interactions? (Link)
@@ A.4 Summarization into Key Points
While I believe the concerns regarding children’s screen time are valid, I believe it is somewhat biased to not take this problem, which is a genuine issue right now, as an everyone problem, and not a student one. With that in mind, I do believe that for the majority of our generation it would not be as hard to disconnect with devices after the pandemic.
I know that I, along with many other teenagers, would like nothing more than to go back to school, play sports outside, meet new people and such. I doubt I could say the same for our older generations, because while parents ban teenagers from using their phones, they often spend many more hours on their devices. Cranky grandparents grumble about how ’back in their day they lived without a mobile phone’ usually forget the point of their stories. Back in their day indeed.
We live in a different era, and to think that the same rules should apply to teenagers, especially during a pandemic, isn’t logical.Pinterest, Youtube, Twitter aren’t just places where teens while away their time. They are also places for new ideas, watching college lectures, and political discourse. Limiting screen time, while in the short run is helpful, is doubtful to do anything in the long run. Banning technology is not stopping kids from using it. It’s just stopping them from telling their parents about it.
I would like to end by saying, parents, please trust your teenager just a little bit more, and don’t be too worried about their screen time!
@@ B.3 Significance Testing for Results on Homogenization
Stereotype
0.02*
0.34
0.01*
Stereotype
0.13
0.18
0.81
School
0.03*
0.05*
0.86
School
0.78
0.15
0.17
Mindfulness
0.00*
0.00*
0.03*
Mindfulness
0.00*
0.11
0.64
Athletes
0.24
0.14
0.53
Athletes
0.09
0.05*
0.67
Audiobook
0.73
0.03*
0.25
Audiobook
0.00*
0.06
0.01*
Animal Welfare
0.67
0.13
0.01*
Animal Welfare
0.31
0.02*
0.00*
Extreme Sports
0.03*
0.02*
0.93
Extreme Sports
0.43
0.11
0.38
Screen Time
0.00*
0.09
0.00*
Screen Time
0.00*
0.00*
0.06*
Dating
0.17*
0.00*
0.04*
Dating
0.01*
0.01*
0.41
Citizens
0.51
0.08
0.16
Citizens
0.16
0.11
0.87
@@
Stereotype
0.02*
0.01*
0.00*
Stereotype
0.02*
0.01*
0.03*
School
0.17
0.12
0.71
School
0.14
0.22
0.62
Mindfulness
0.00*
0.01*
0.42
Mindfulness
0.02*
0.22
0.81
Athletes
0.00*
0.00*
0.52
Athletes
0.00*
0.00*
0.90
Audiobook
0.01*
0.15
0.69
Audiobook
0.00*
0.00*
0.76
Animal Welfare
0.49
0.11
0.15
Animal Welfare
0.13
0.00*
0.00*
Extreme Sports
0.35
0.00*
0.00*
Extreme Sports
0.04*
0.00*
0.09
Screen Time
0.00*
0.00*
0.00*
Screen Time
0.00*
0.00*
0.70
Dating
0.00*
0.61
0.00*
Dating
0.88
0.01*
0.02*
Citizens
0.00*
0.14
0.11
Citizens
0.00*
0.04*
0.11
@@ Appendix C Additional Results
Key point Level Rouge-L
0.1536
0.1578
0.1660
BertScore
0.6305
0.6284
0.6437
Essay Level Rouge-L
0.1498
0.1523
0.1621
BertScore
0.6044
0.6130
0.6318
@@
LZMA
0.305
0.302
0.293
ZLIB
0.353
0.351
0.342
GZIP
0.352
0.341
0.350

### v3

@@ 3 How much do users engage with the model?
While I believe the concerns regarding children’s screen time are valid, I believe it is somewhat biased to not take this problem, which is a genuine issue right now, as an everyone problem, [skipped] I know that I, along with many other teenagers, would like nothing more than to go back to school, play sports outside, meet new people, and such. [skipped] They are also places for new ideas, watching college lectures, and political discourse [skipped]
  • The problem of screen time should be considered an everyone problem, not just a student one
  • Social media can be used for educational purposes
  • Limiting screen time may not be effective in the long run
  • Parents should trust their teenager more and not worry too much about their screen time
@@ 5.2 Results
keeping up with the news 7 keep up with the news 14
in my opinion the most 7 on animal welfare when humans 12
keep up with the news 6 to focus on animal welfare 11
opinion the most important things 6 selfish to pursue risky sports 11
but on the other hand 5 students should learn in school 11
the most important thing that 5 wrong to focus on animal 10
wrong to focus on animal 5 sports like extreme mountain climbing 10
focus on animal welfare when 5 keeping up with the news 9
unfair when it is considered 5 the end of the day 9
in my opinion listening to 4 things students should learn in 9
@@ 7 Related Work
Traditional evaluation metrics primarily assess text similarity to reference text (Papineni et al., 2002; Lin, 2004; Sellam et al., 2020). Pillutla et al. (2021); Liu et al. (2021); Pillutla et al. (2023) propose Mauve, a metric that measures the match between the distribution of output produced by a generative model and a reference distribution. Their experiments show a trade-off between the quality and diversity of output generated by both image and language models, investigating different model sizes and decoding strategies. Our work builds on the recommendations in Pillutla et al. (2023) to propose metrics for diversity that can be used across model families and scales, further showing the relationship between diversity and aligning LLMs with human feedback. In interactive contexts where references are absent, alternative metrics gauge model assistance effectiveness indirectly. These include user-rated model helpfulness (Roemmele and Gordon, 2018; Clark et al., 2018; Padmakumar and He, 2022), retention of model suggestions in the final text (Akoury et al., 2020), and completion time (Buschek et al., 2021). Lee et al. (2022b) additionally considers user ownership and enjoyment which emphasizes maintaining user writing style. Our work extends existing metrics that evaluate lexical diversity and text style based on token and POS n-grams statistics (Roemmele et al., 2017; See et al., 2019; Tevet and Berant, 2021; Meister et al., 2023). We adapt these metrics for co-writing scenarios and introduce content-specific diversity measures.
@@ Acknowledgements
We would like to thank Mina Lee, Nick Lourie, Naomi Saphra, Richard Pang, and Nitish Joshi for their input at various stages of the project. We would also like to thank Abby Rabinowitz and Alexander Landfair from the NYU Expository Writing Program for valuable discussions during the planning stages. We would finally like to acknowledge all the writers recruited for this project without whom this work would not have been possible. This work is supported by Open Philanthropy, AWS AI, and the National Science Foundation under Grant No. 1922658 and Grant No. IIS-2340345.
@@ A.2 Decoding Parameters on OpenAI API
school What Are the Most Important Things Students Should Learn in School? In your opinion, what are the most important things students should learn in school? What is the most important thing you have learned in school? How has this knowledge affected your life? How do you think it will help your success in the future? (Link)
@@
stereotype What Stereotypical Characters Make You Cringe? What stereotypical characters in books, movies or television shows make you cringe and why? Would you ever not watch or read something because of its offensive portrayal of someone? (Link)
@@
audiobook Is Listening to a Book Just as Good as Reading It? Do you listen to audiobooks? What are the benefits, in your opinion, of listening instead of reading? Are there advantages to reading that cannot be gained by listening? Which method do you prefer? Why? (Link)
@@
athletes Should College Athletes Be Paid? Do you think college athletes should be paid? Or is a college scholarship and other non-monetary perks like the opportunity to play in front of cheering fans enough? What possible difficulties or downsides might there be in providing monetary compensation to players? (Link)
@@
extremesports Is It Selfish to Pursue Risky Sports Like Extreme Mountain Climbing? Some sports, like extreme mountain climbing, are dangerous. Since there are varying degrees of risk in most, if not all, sports (such as the possibility of concussions, broken bones and even death), how does one decide where the line might be drawn between what is reasonable and what is not? Are some sports simply too dangerous to be called a sport? (Link)
@@
animal Is It Wrong to Focus on Animal Welfare When Humans Are Suffering? Would you be surprised to hear that a study found that research subjects were more upset by stories of a dog beaten by a baseball bat than of an adult similarly beaten? Or that other researchers found that if forced to choose, 40 percent of people would save their pet dog over a foreign tourist. Why do you think many people are more empathetic toward the suffering of animals than that of people? In your opinion, is it wrong to focus on animal welfare when humans are suffering? Why do you think so? (Link)
@@
news Are We Being Bad Citizens If We Don’t Keep Up With the News? In your opinion, are we being bad citizens if we don’t keep up with the news? Do you think all people have some responsibility to know what is going on in the world? Does engaging with current events actually do anything at all? Why do you think the way you do? (Link)
@@
mindfulness Should Schools Teach Mindfulness? Have you ever tried mindfulness or meditation, practices that focus on the present moment and being aware of your thoughts, feelings and bodily sensations? If so, what was it like for you? If not, does it sound like something you’d like to try? Do you think that such practices have a place in schools? Why or why not? (Link)
@@
screen How Worried Should We Be About Screen Time During the Pandemic? The coronavirus pandemic ended the screen time debate: Screens won. We all now find ourselves on our screens for school, for work and for connecting with family and friends during this time of social distancing and increased isolation. But should we be worried about this excessive screen use right now? Or should we finally get over it and embrace the benefits of our digital devices? (Link)
@@
dating How Do You Think Technology Affects Dating? Have you had any experience with dating? Have you ever used dating apps? If so, what has it been like for you? If not, why not? How do you think technology — like apps, Netflix, social media and texting — affects dating and relationships? In your opinion, does it improve or worsen romantic interactions? (Link)
@@ A.4 Summarization into Key Points
While I believe the concerns regarding children’s screen time are valid, I believe it is somewhat biased to not take this problem, which is a genuine issue right now, as an everyone problem, and not a student one. With that in mind, I do believe that for the majority of our generation it would not be as hard to disconnect with devices after the pandemic. I know that I, along with many other teenagers, would like nothing more than to go back to school, play sports outside, meet new people and such. I doubt I could say the same for our older generations, because while parents ban teenagers from using their phones, they often spend many more hours on their devices. Cranky grandparents grumble about how ’back in their day they lived without a mobile phone’ usually forget the point of their stories. Back in their day indeed. We live in a different era, and to think that the same rules should apply to teenagers, especially during a pandemic, isn’t logical.Pinterest, Youtube, Twitter aren’t just places where teens while away their time. They are also places for new ideas, watching college lectures, and political discourse. Limiting screen time, while in the short run is helpful, is doubtful to do anything in the long run. Banning technology is not stopping kids from using it. It’s just stopping them from telling their parents about it. I would like to end by saying, parents, please trust your teenager just a little bit more, and don’t be too worried about their screen time!
@@ B.3 Significance Testing for Results on Homogenization
Stereotype 0.02* 0.34 0.01* Stereotype 0.13 0.18 0.81
School 0.03* 0.05* 0.86 School 0.78 0.15 0.17
Mindfulness 0.00* 0.00* 0.03* Mindfulness 0.00* 0.11 0.64
Athletes 0.24 0.14 0.53 Athletes 0.09 0.05* 0.67
Audiobook 0.73 0.03* 0.25 Audiobook 0.00* 0.06 0.01*
Animal Welfare 0.67 0.13 0.01* Animal Welfare 0.31 0.02* 0.00*
Extreme Sports 0.03* 0.02* 0.93 Extreme Sports 0.43 0.11 0.38
Screen Time 0.00* 0.09 0.00* Screen Time 0.00* 0.00* 0.06*
Dating 0.17* 0.00* 0.04* Dating 0.01* 0.01* 0.41
Citizens 0.51 0.08 0.16 Citizens 0.16 0.11 0.87
@@
Stereotype 0.02* 0.01* 0.00* Stereotype 0.02* 0.01* 0.03*
School 0.17 0.12 0.71 School 0.14 0.22 0.62
Mindfulness 0.00* 0.01* 0.42 Mindfulness 0.02* 0.22 0.81
Athletes 0.00* 0.00* 0.52 Athletes 0.00* 0.00* 0.90
Audiobook 0.01* 0.15 0.69 Audiobook 0.00* 0.00* 0.76
Animal Welfare 0.49 0.11 0.15 Animal Welfare 0.13 0.00* 0.00*
Extreme Sports 0.35 0.00* 0.00* Extreme Sports 0.04* 0.00* 0.09
Screen Time 0.00* 0.00* 0.00* Screen Time 0.00* 0.00* 0.70
Dating 0.00* 0.61 0.00* Dating 0.88 0.01* 0.02*
Citizens 0.00* 0.14 0.11 Citizens 0.00* 0.04* 0.11
@@ Appendix C Additional Results
Key point Level Rouge-L 0.1536 0.1578 0.1660
BertScore 0.6305 0.6284 0.6437
Essay Level Rouge-L 0.1498 0.1523 0.1621
BertScore 0.6044 0.6130 0.6318
@@
LZMA 0.305 0.302 0.293
ZLIB 0.353 0.351 0.342
GZIP 0.352 0.341 0.350